\clearpage
\begin{thebibliography}{00}
% \bibitem{label}
% Text of bibliographic item
% notes:
% \bibitem{label} \note
% subbibitems:
% \begin{subbibitems}{label}
% \bibitem{label1}
% \bibitem{label2}
% If there is a note, it should come last:
% \bibitem{label3} \note
% \end{subbibitems}
\bibitem{AD09} A.~S. Adekola, Ph. D. thesis, Ohio University (2009).
\bibitem{AI70} J.~H. Aitken et al., Can. J. Phys. 48 (1970) 1617.
\bibitem{Ajz72} F. Ajzenberg-Selove, Nucl. Phys. A 190 (1972) 1.
\bibitem{Ajz91} F. Ajzenberg-Selove, Nucl. Phys. A 523 (1991) 1.
\bibitem{AL75} R. Almanza et al., Nucl. Phys. A 248 (1975) 214.
\bibitem{raman}N. Anantaraman et al., Nucl. Phys. A 279 (1977) 474.
\bibitem{Ang99} C. Angulo et al., Nucl. Phys. A 656 (1999) 3.
\bibitem{antilla} A. Anttila, J. Keinonen and M. Bister, J. Phys. G 3 (1977) 1241.
\bibitem{Ara06} A. Arazi et al., Phys. Rev. C 74 (2006) 025802.
\bibitem{Arn06} M. Arnould and S. Goriely, Nucl. Phys. A 777 (2006) 157.
\bibitem{Aud03} G. Audi, A.~H. Wapstra and C. Thibault, Nucl. Phys. A 729 (2003) 337.
\bibitem{Axe98} L. Axelsson et al., Nucl. Phys. A 634 (1998) 475.
\bibitem{Bar00} D.~W. Bardayan et al., Phys. Rev. C 62 (2000) 055804.
\bibitem{BA01} D.~W.Bardayan et al., Phys. Rev. C 63 (2001) 065802.
\bibitem{BA02} D.~W. Bardayan et al., Phys. Rev. Lett. 89 (2002) 262501.
\bibitem{Bar02} D.~W. Bardayan et al., Phys. Rev. C 65 (2002) 032801(R).
\bibitem{BA04} D.~W. Bardayan et al., Phys. Rev. C 70 (2004) 015804.
\bibitem{BA05} D.~W. Bardayan, R.~L. Kozub and M.~S. Smith, Phys. Rev. C 71 (2005) 018801.
\bibitem{Bar06} D.~W. Bardayan et al., Phys. Rev. C 74 (2006) 045804.
\bibitem{Bar07} D.~W. Bardayan et al., Phys. Rev. C 76 (2007) 045803.
\bibitem{BA55} G.~A. Bartholomew et al., Can. J. Phys. 33 (1955) 441.
\bibitem{Bat01} N. Bateman et al., Phys. Rev. C 63 (2001) 035803.
\bibitem{Bax73} A.~M. Baxter and S. Hinds, Nucl. Phys. A 211 (1973) 7.
\bibitem{BE82} H.~W. Becker et al., Z. Phys. A 305 (1982) 319.
\bibitem{BE95} H.~W. Becker et al., Z. Phys. A 351 (1995) 453.
\bibitem{Ben76} W. Benenson et al., Phys. Rev. C 13 (1976) 1479.
\bibitem{Ben77} W. Benenson et al., Phys. Rev. C 15 (1977) 1187.
\bibitem{Ber79} U.~E.~P. Berg and K. Wiehard, Nucl. Phys. A 318 (1979) 453.
\bibitem{BO75} J. Bommer et al., Nucl. Phys. A 251 (1975) 246.
\bibitem{Bos68} B. Bosnjakovic et al., Nucl. Phys. A 110 (1968) 17.
\bibitem{Buc88} L. Buchmann, J.~M. D'Auria and P. McCorquodale, Astrophys. J. 324 (1988) 953.
\bibitem{Cag01} J.~A. Caggiano et al., Phys. Rev. C 64 (2001) 025802.
\bibitem{Cag02} J.~A. Caggiano et al., Phys. Rev. C 65 (2002) 055801.
\bibitem{CH06} K.~Y. Chae et al., Phys. Rev. C 74 (2006) 012801(R).  
\bibitem{Cha07} A. Chafa et al., Phys. Rev. C 75 (2007) 035810.
\bibitem{CH86} A.~E. Champagne and M. Pitt, Nucl. Phys. A 457 (1986) 367.
\bibitem{Cha88} A.~E. Champagne et al., Nucl. Phys. A 487 (1988) 433.
\bibitem{Cha90} A.~E. Champagne, P.~V. Magnus and M.~S. Smith, Nucl. Phys. A 512 (1990) 317.
\bibitem{Cha93} A.~E. Champagne, B.~A. Brown and R. Sherr, Nucl. Phys. A 556 (1993) 123.
\bibitem{Cha06} R. Chatterjee, A. Okolowicz and M. Ploszajczak, Nucl. Phys. A 764 (2006) 528.
\bibitem{Cho70} G. Chouraqui et al., J. de Phys. 31 (1970) 249.
\bibitem{Chr99} C. Chronidou et al., Eur. Phys. J. A 6 (1999) 303.
\bibitem{Cle04} R.~R.~C. Clement et al., Phys. Rev. Lett. 92 (2004) 172502. 
\bibitem{Com07} H. Comisel et al., Phys. Rev. C 75 (2007) 045807.
\bibitem{CO94} R. Coszach et al., Phys. Rev. C 50 (1994) 1695.
\bibitem{CO04} M. Couder et al., Phys. Rev. C 69 (2004) 022801.
\bibitem{Cow98} G. Cowan, {\it Statistical Data Analysis} (Oxford University Press, New York, 1998).
\bibitem{cseh} J. Cseh et al., Nucl. Phys. A 385 (1982) 43.
\bibitem{CU81} A. Cunsolo et al., Phys.Rev. C 24 (1981) 476. 
\bibitem{Dab03} Dababneh et al., Phys. Rev. C 68 (2003) 025801.
\bibitem{Dal09} J.~C. Dalouzy et al., Phys. Rev. C, in print (2009).
\bibitem{DAu04} J.~M. D'Auria et al., Phys. Rev. C 69 (2004) 065804.
\bibitem{Dav03b} B. Davids et al., Phys. Rev. C 67 (2003) 065808.
\bibitem{OL96} F. de Oliveira et al., Nucl. Phys. A 587 (1996) 231; F. de Oliveira, thesis, Universit\'e Paris Sud, unpublished (1995).
\bibitem{OL97} F. de Oliveira et al., Phys. Rev. C 55 (1997) 3149.
\bibitem{Dei08} C.~M. Deibel, Ph. D. thesis, Yale University (2008).
\bibitem{Des88} P. Descouvemont, Phys. Rev. C 38 (1988) 2397.
\bibitem{Des00} P. Descouvemont, Astrophys. J. 543 (2000) 425.
\bibitem{Des04} P. Descouvemont et al., At. Data Nucl. Data Tab. 88 (2004) 203. 
\bibitem{DI71a} W.~R. Dixon and R.~S. Storey, Can. J. Phys. 49 (1971) 1714.
\bibitem{DI71b} W.~R. Dixon et al., Phys. Rev. Lett. 27 (1971) 1460.
\bibitem{DI77} W.~R. Dixon and R.~S. Storey, Nucl. Phys. A 284 (1977) 97.
\bibitem{Doo07} P. Doornenbal et al., Phys. Lett. B 647 (2007) 237.
\bibitem{dr74} J.~P. Draayer et al., Phys. Lett. B 53 (1974) 250. 
\bibitem{Duf00} M. Dufour and P. Descouvemont, Nucl. Phys. A 672 (2000) 153.
\bibitem{Duf04} M. Dufour and P. Descouvemont, Nucl. Phys. A 730 (2004) 316.
\bibitem{DU07} M. Dufour and P. Descouvemont, Nucl. Phys. A 785 (2007) 381.
\bibitem{End77} P.~M. Endt, At. Data Nucl. Data Tab. 19 (1977) 23.
\bibitem{End90} P.~M. Endt, Nucl. Phys. A 521 (1990) 1.
\bibitem{End98} P.~M. Endt, Nucl. Phys. A 633 (1998) 1.
\bibitem{EnB93} P.~M. Endt and J.~G.~L. Booten, Nucl. Phys. A 555 (1993) 499.
\bibitem{End78} P.~M. Endt and C. van der Leun, Nucl. Phys. A 310 (1978) 1.
\bibitem{EnR87} P.~M. Endt and C. Rolfs, Nucl. Phys. A 467 (1987) 261.
\bibitem{Eng05} S. Engel et al., Nucl. Instr. Meth. A 553 (2005) 491.
\bibitem{FE59} A.~J. Ferguson and H.~E. Gove, Can. J. Phys. 37 (1959) 660.
\bibitem{Fin74} C.~L. Fink and J.~P. Schiffer, Nucl. Phys. A 225 (1974) 93.
\bibitem{endsf} R.~B. Firestone, Nucl. Data Sheets 103 (2004) 269.
\bibitem{endsf2} R.~B. Firestone, Nucl. Data Sheets 108 (2007) 2319. 
\bibitem{fif} L.~K. Fifield et al., Nucl. Phys. A 309 (1978) 77; Nucl. Phys. A 322 (1979) 1. 
\bibitem{For08} A. Formicola et al., J. Phys. G 35 (2008) 014013.
\bibitem{FO00} H.~T. Fortune, R. Sherr and B.~A. Brown, Phys. Rev. C 61 (2000) 057303.
\bibitem{Fox05} C. Fox et al., Phys. Rev. C 71 (2005) 055801.
\bibitem{FR61} J.~B. French, S. Iwao and E. Vogt, Phys. Rev. 122 (1961) 1248.
\bibitem{Fyn00} H.~O.~U. Fynbo et al., Nucl. Phys. A 677 (2000) 38.
\bibitem{Gad08} A. Gade et al., Phys. Rev. C 77 (2008) 044306.
\bibitem{GA87} M. Gai et al., Phys. Rev. C 36 (1987) 1256.
\bibitem{Gie94} U. Giesen et al., Nucl. Phys. A 567 (1994) 146.
\bibitem{Goe00} J. G\"orres et al., Phys. Rev. C 62, 055801 (2000).
\bibitem{Gol04} V.~Z. Goldberg et al., Phys. Rev. C 69 (2004) 024602.
\bibitem{Gom05} T. Gomi et al., J. Phys. G 31 (2005) S1517.
\bibitem{Gor83} J. G\"orres et al., Nucl. Phys. A 408 (1983) 372.
\bibitem{GO90} J. G\"orres et al., Nucl. Phys. A 517 (1990) 329.
\bibitem{GO92} J. G\"orres et al., Nucl. Phys. A 548 (1992) 414.
\bibitem{Gor08} S. Goriely, S. Hilaire and A.~J. Koning, Astron. Astrophys. 487 (2008) 767, and {\it private communication}.
\bibitem{Gra68} P.~J. Graff et al., J. de Phys. 29 (1968) 141.
\bibitem{Gra90} S. Graff et al., Nucl. Phys. A 510 (1990) 346.
\bibitem{Gre91} M.~B. Greenfield et al., Nucl. Phys. A 524 (1991) 228.
\bibitem{Guo06} B. Guo et al., Phys. Rev. C 73 (2006) 048801.
\bibitem{Hah96} K.~I. Hahn et al., Phys. Rev. C 54 (1996) 1999.
\bibitem{Hal01} S.~E. Hale et al., Phys. Rev. C 65 (2001) 015801.
\bibitem{Hal04} S.~E. Hale et al., Phys. Rev. C 70 (2004) 045802.
\bibitem{Han00} V.~Y. Hansper et al., Phys. Rev. C 61 (2000) 028801.
\bibitem{Har00} S. Harissopulos et al., Eur. Phys. J. A 9 (2000) 479.
\bibitem{He07} J.~J. He et al., Phys. Rev. C 76 (2007) 055802.
\bibitem{Her95} H. Herndl et al., Phys. Rev. C 52 (1995) 1078.
\bibitem{Her98} H. Herndl et al., Phys. Rev. C 58 (1998) 1798.
\bibitem{high} G.~J. Highland and T.~T. Thwaites, Nucl. Phys. A 109 (1968) 163.
\bibitem{Hof99} R.~D. Hoffman et al., Astrophys. J. 521 (1999) 735.
\bibitem{How70} A.~J. Howard et al., Nucl. Phys. A 152 (1970) 317.
\bibitem{Ili89} C. Iliadis, Diplom thesis (Univ. of M\"unster, 1989).
\bibitem{Ili90} C. Iliadis et al., Nucl. Phys. A 512 (1990) 509.
\bibitem{Ili93} C. Iliadis et al., Nucl. Phys. A 559 (1993) 83.
\bibitem{Ili94} C. Iliadis et al., Nucl. Phys. A 571 (1994) 132.
\bibitem{Ili96} C. Iliadis et al., Phys. Rev. C 53 (1996) 475. 
\bibitem{ili2} C. Iliadis, Nucl. Phys. A 618 (1997) 166. 
\bibitem{Ili99} C. Iliadis et al., Astrophys. J. 524 (1999) 434.
\bibitem{Ili01} C. Iliadis et al., Astrophys. J. Suppl. 134 (2001) 151.
\bibitem{Ili04} C. Iliadis and M. Wiescher, Phys. Rev. C 69 (2004) 064305.
\bibitem{Ili07} C. Iliadis, {\it Nuclear Physics of Stars} (Wiley-VCH, Weinheim, 2007).
\bibitem{Ili08} C. Iliadis et al., Phys. Rev. C 77 (2008) 045802.
\bibitem{Jen04} D.~G. Jenkins et al., Phys. Rev. Lett. 92 (2004) 031101.
\bibitem{Jen06} D.~G. Jenkins et al., Phys. Rev. C 73 (2006) 065802.
\bibitem{Joh74} P.~M. Johnson, M.~A. Meyer and D. Reitmann, Nucl. Phys. A 218 (1974) 333.
\bibitem{Kal78} J. Kalifa et al., Phys. Rev. C 17 (1978) 1961.
\bibitem{Kan06} R. Kanungo et al., Phys. Rev. C 74 (2006) 045803.
\bibitem{Kei77} J. Keinonen, M. Riihonen and A. Anttila, Phys. Rev. C 15 (1977) 579.
\bibitem{Kie79} W.~E. Kieser, R.~E. Azuma and K.~P. Jackson, Nucl. Phys. A 331 (1979) 155. 
\bibitem{KU71} H.~M. Kuan, C.~J. Umbarger and D.~G. Shirk, Nucl. Phys. A 160 (1971) 211; Erratum: Nucl. Phys. A 196 (1972) 634.
\bibitem{KU76} H.~M. Kuan and D.~G. Shirk, Phys. Rev. C 13 (1976) 883.
\bibitem{Kub92} S. Kubono et al., Nucl. Phys. A 537 (1992) 153.
\bibitem{Kub95} S. Kubono, T. Kajino and S. Kato, Nucl. Phys. A 588 (1995) 521.
\bibitem{La08} M. La Cognata et al., Phys. Rev. Lett. 101 (2008) 152501.
\bibitem{Lan86} K.~H. Langanke et al., Astrophys. J. 301 (1986) 629.
\bibitem{Li76} T.~K. Li et al., Phys. Rev. C 13 (1976) 55.
\bibitem{Lon09} R. Longland et al., in print (2009).
\bibitem{LO79} H. Lorentz-Wirzba et al., Nucl. Phys. A 313 (1979) 346. 
\bibitem{Lot08} G. Lotay et al., Phys. Rev. C 77 (2008) 042802(R).
\bibitem{Lot09} G. Lotay et al., Phys. Rev. Lett. 102 (2009) 162502.
\bibitem{LU04} M. Lugaro et al., Astrophys. J. 615 (2004) 934.
\bibitem{ly69} P.~B. Lyons et al., Nucl. Phys. A 130 (1969) 25.
\bibitem{Ma07} Z. Ma et al., Phys. Rev. C 76 (2007) 015803.
\bibitem{maas} J.~W. Maas et al., Nucl. Phys. A 301 (1978) 213.
\bibitem{Mac73} H. Mackh et al., Nucl. Phys. A 202 (1973) 497.
\bibitem{Mac80} J.~D. MacArthur et al., Phys. Rev. C 22 (1980) 356.
\bibitem{Mac85} J.~D. MacArthur et al., Phys. Rev. C 32 (1985) 314.
\bibitem{MA87} P.~V. Magnus et al., Nucl. Phys. A 470 (1987) 206.
\bibitem{MA78} H.~B. Mak et al., Nucl. Phys. A 304 (1978) 210.
\bibitem{Mao95} Z.~Q. Mao, H.~T. Fortune and A.~G. Lacaze, Phys. Rev. Lett. 74 (1995) 3760.
\bibitem{Mao96} Z.~Q. Mao, H.~T. Fortune and A.~G. Lacaze, Phys. Rev. C 53 (1996) 1197.
\bibitem{May01} A. Mayer, Ph. D. thesis (Universit\"at Stuttgart, 2001).
\bibitem{Moa07} B.~H. Moazen et al., Phys Rev. C 75 (2007) 065801.
\bibitem{Moh05} P. Mohr, Phys. Rev. C 72 (2005) 035803.
\bibitem{Moo05} J.~Y. Moon et al., Nucl. Phys. A 758 (2005) 158c.
\bibitem{Muk04} M. Mukherjee et al., Phys. Rev. Lett., 93 (2004) 150801.
\bibitem{muk2} M. Mukherjee et al., Eur. Phys. J. A 35 (2008) 37.
\bibitem{MU79} G. Murillo et al., Nucl. Phys. A 318 (1979) 352.
\bibitem{Mur06} A.~St.~J. Murphy et al., Phys. Rev. C 73 (2006) 034320.
\bibitem{Myt08} S.~Mythili et al., Phys. Rev. C 77 (2008) 035803.
\bibitem{NE07} C.~D. Nesaraja et al., Phys. Rev. C 75 (2007) 055809.
\bibitem{New07} J.~R. Newton et al., Phys. Rev. C 75 (2007) 055808.
\bibitem{New08} J.~R. Newton, R. Longland and C. Iliadis, Phys. Rev. C 78 (2008) 025805.
\bibitem{NI77} M. Niecke et al., Nucl. Phys. A 289 (1977) 408.
\bibitem{OR73} H. Orihara, G. Rudolf and Ph. Gorodetzky, Nucl. Phys. A 203 (1973) 78.
\bibitem{Par05} A. Parikh et al., Phys. Rev. C 71 (2005) 055804.
\bibitem{Par99} S.~H. Park et al., Phys. Rev. C 59 (1999) 1182. 
\bibitem{Par04} Y. Parpottas et al., Phys. Rev. C 70 (2004) 065805; and Phys. Rev. C 73 (2006) 049907(E).
\bibitem{Pow71} J.~R. Powers et al., Phys. Rev. C 4 (1971) 2030.
\bibitem{Pow99} D.~C. Powell et al., Nucl. Phys. A 660 (1999) 349.
\bibitem{Pre92} W.~H. Press, S.~A. Teukolsky, W.~T. Vetterling and B.~P. Flannery, Numerical Recipes (Cambridge University Press, Cambridge, 1992).
\bibitem{PR89} D.~M. Pringle and W.~J. Vermeer, Nucl. Phys. A 499 (1989) 117.
\bibitem{RA72} J.~J. Ramirez, R.~A. Blue and H.~R. Weller, Phys. Rev. C 5 (1972) 17.
\bibitem{Rau00} T. Rauscher and F.-K. Thielemann, At. Data Nucl. Data Tab. 75 (2000) 1.
\bibitem{Roe02} H. R\"opke, J. Brenneisen and M. Lickert, Eur. J. A 14 (2002) 159.
\bibitem{RO72a} D.~W.~O. Rogers, J.~H. Aitken and A.~E. Litherland, Can. J. Phys. 50 (1972) 268.
\bibitem{RO72b} D.~W.~O. Rogers, R.~P. Beukens and W.~T. Diamond, Can. J. Phys. 50 (1972) 2428.
\bibitem{RO76} D.~W.~O. Rogers et al., Can. J. Phys. 54 (1976) 938.
\bibitem{Rol75} C. Rolfs et al., Nucl. Phys. A 241 (1975) 460.
\bibitem{Rol73a} C. Rolfs, I. Berka and R.~E. Azuma, Nucl. Phys. A 199 (1973) 306.
\bibitem{Rol73b} C. Rolfs, A.~M. Charlesworth and R.~E. Azuma, Nucl. Phys. A 199 (1973) 257.
\bibitem{Ros95} J.~G. Ross et al., Phys. Rev. C 52 (1995) 1681.
\bibitem{Row02} C. Rowland et al., Phys. Rev. C 65 (2002) 064609.
\bibitem{Row04} C. Rowland et al., Astrophys. J. 615 (2004) L37.
\bibitem{Rui05} C. Ruiz et al., Phys. Rev. C 71 (2005) 025802.
\bibitem{Rui06} C. Ruiz et al., Phys. Rev. Lett. 96 (2006) 252501.
\bibitem{sch} P. Schmalbrock et al., Nucl. Phys. A 398 (1983) 279.
\bibitem{Sch95} S. Schmidt et al., Nucl. Phys. A 591 (1995) 227.
\bibitem{Sch97} H. Schatz et al., Phys. Rev. Lett. 79 (1997) 3845.
\bibitem{Sch05} H. Schatz et al., Phys. Rev. C 72 (2005) 065804.
\bibitem{SE69} D.~L. Sellin, H.~W. Newson and E.~G. Bilpuch, Ann. of Phys. 51 (1969) 461.
\bibitem{Seu90} S. Seuthe et al., Nucl. Phys. A 514 (1990) 471.
\bibitem{Sew05} D. Seweryniak et al., Phys. Rev. Lett. 94 (2005) 032501.
\bibitem{Sew07} D. Seweryniak et al., Phys. Rev. C 75 (2007) 062801(R).
\bibitem{smu} P.~J.~M. Smulders, Physica 31 (1965) 973.
\bibitem{smE} P.~J.~M. Smulders and P.~M. Endt, Physica 28 (1962) 1093.
\bibitem{Ste74} M.~A. Stephens, J. Amer. Stat. Assoc. 69 (1974) 730.
\bibitem{Ste96} F. Stegm\"uller et al., Nucl. Phys. A 601 (1996) 168.
\bibitem{str08} E. Strandberg et al., Phys. Rev. C 77 (2008) 055801.
\bibitem{SY78} T.~J. Symons et al., J. Phys. G 4 (1978) 411.
\bibitem{Tan05} W.~P. Tan et al., Phys. Rev. C 72 (2005) 041302.
\bibitem{Tan81} T. Tanabe et al., Phys. Rev. C 6 (1981) 2556.
\bibitem{ta83} T. Tanabe et al., Nucl. Phys. A 399 (1983) 241.
\bibitem{Tho99} W.~J. Thompson and C. Iliadis, Nucl. Phys. A 647 (1999) 259.
\bibitem{Til93} D.~R. Tilley et al., Nucl. Phys. A 564 (1993) 1.
\bibitem{Til95} D.~R. Tilley et al., Nucl. Phys. A, 595 (1995) 1.
\bibitem{Til98} D.~R. Tilley et al., Nucl. Phys. A 636 (1998) 249.
\bibitem{Tim88} R. Timmermann et al., Nucl. Phys. A 477 (1988) 105.
\bibitem{Tom04} I. Tomandl et al., Phys. Rev. C 69 (2004) 014312.
\bibitem{Tra75} H.~P. Trautvetter, Nucl. Phys. A 243 (1975) 37.
\bibitem{Tra78} H.~P. Trautvetter et al., Nucl. Phys. A 297 (1978) 489.
\bibitem{Tri99} W. Trinder et al., Phys. Lett. B 459 (1999) 67.
\bibitem{UN74} B.~Y. Underwood et al., Nucl. Phys. A 225 (1974) 253.
\bibitem{UT98} S. Utku et al., Phys. Rev. C 57 (1998) 2731.  
\bibitem{VA98} G.~Vancraeynest et al., Phys. Rev. C 57 (1998) 2711.
\bibitem{Vis07} D.~W. Visser et al., Phys. Rev. C 76 (2007) 065803.
\bibitem{Vis08} D.~W. Visser et al., Phys. Rev. C 78 (2008) 028802.
\bibitem{Vog89} R.~B. Vogelaar, Ph.D. thesis (Caltech, 1989).
\bibitem{VO90} R.~B. Vogelaar et al., Phys. Rev. C 42 (1990) 753.
\bibitem{Vog96} R.~B. Vogelaar et al., Phys. Rev. C 53 (1996) 1945.
\bibitem{WI80} M. Wiescher et al., Nucl. Phys. A 349 (1980) 165.
\bibitem{WI02} S. Wilmes et al., Phys. Rev. C 66 (2002) 065802.
\bibitem{Wre07} C. Wrede et al., Phys. Rev. C 76 (2007) 052802(R).
\bibitem{Wre09} C. Wrede et al., Phys. Rev. C 79 (2009) 045808.
\bibitem{Wre09b} C. Wrede, Phys. Rev. C 79 (2009) 035803.
\bibitem{YA62} K. Yagi, J. Phys. Soc. Japan 17 (1962) 604.
\bibitem{Yaz07} C. Yazidjian et al., Phys. Rev. C 76 (2007) 024308.
\bibitem{Yok82} H. Yokota et al., Nucl. Phys. A 383 (1982) 298.
\bibitem{Yon06} K. Yoneda et al., Phys. Rev. C 74 (2006) 021303(R).
\end{thebibliography}

\documentclass{elsart1p}

\usepackage{graphicx}

\usepackage{amssymb,longtable,amsmath}

\begin{document}

\begin{frontmatter}



\title{Charged-Particle Thermonuclear Reaction Rates:\\ III. Nuclear Physics Input}


\author{C. Iliadis}, \author{R. Longland}, \author{A. E. Champagne}
\address{Department of Physics and Astronomy, University of North Carolina, Chapel Hill, NC 27599-3255, USA; Triangle Universities Nuclear Laboratory, Durham, NC 27708-0308, USA}
\author{A. Coc}
\address{Centre de Spectrom\'etrie Nucl\'eaire et de Spectrom\'etrie de Masse (CSNSM), UMR 8609, CNRS/IN2P3 and Universit\'e
        Paris Sud 11, B\^atiment 104, 91405 Orsay Campus, France}

\begin{abstract}
The nuclear physics input used to compute the Monte Carlo reaction rates and probability density functions that are tabulated in the second paper of this series (Paper II) is presented. Specifically, we publish the input files to the Monte Carlo reaction rate code \texttt{RatesMC}, which is based on the formalism presented in the first paper of this series  (Paper I). This data base contains overwhelmingly experimental nuclear physics information. The survey of literature for this review was concluded in November 2009. 
\end{abstract}


\end{frontmatter}

\section{Introduction}\label{intro}
For a number of reasons, the present evaluation of charged-particle thermonuclear reaction rates represents a significant step forward compared to previous work. First, we developed a new method of computing reaction rates, which is based on Monte Carlo techniques and assigns to each nuclear input quantity a physically motivated probability density function. The method is described in the first paper of this series (Paper I) and the numerical results for reaction rates and rate probability density functions are presented in the second paper of this series (Paper II). Second, a number of years have passed since the last two evaluations \cite{Ang99,Ili01} of similar scope as the present work have been published. The rapid progress seen in the field of nuclear astrophysics over the past few years clearly warrants a new charged-particle reaction rate evaluation. Third, thermonuclear reaction rates are not directly measured quantities, but are derived from a multitude of different nuclear physics input quantities. Consequently, the quality and reliability of a reaction rate evaluation hinges directly on the transparency and reproducibility of the input data. The last two aspects will be addressed in the present paper.

We present here the input files to the Monte Carlo code \texttt{RatesMC} containing the nuclear physics data used to compute our reaction rates. For each reaction we list nuclear properties; nonresonant S-factors; recommended resonance energies, strengths and partial widths; upper limits of resonance contributions; numerical rate integrations; and interferences between resonances of same spin and parity. Section 2 contains a brief discussion of general procedures, literature sources and the status of data. In Sec. 3 we focus attention on a number of important issues. A brief summary is given in Sec. 4. The meaning of each input row in a sample input file to \texttt{RatesMC} is explained in App. \ref{sec:inptfiles}. The actual nuclear physics input data for each of the reactions evaluated in the present work are listed in App. \ref{sec:nuclinput}. 

\section{Procedures, sources and status of data}

We will briefly describe our procedures for nuclear data analysis and evaluation. The details are too numerous to list here and depend on a case-by-case basis, but some general principles can be outlined. All the available sources of literature have been consulted. We focus our attention on refereed journals, but in exceptional circumstances conference proceedings and Ph.D. theses are also taken into account. If a particular quantity has been measured more than once, we adopt a weighted average (except for resonance strengths; see below), unless there was reason to exclude unreliable measurements. In some cases we succeeded in correcting original data for systematic effects, for example, improved stoichiometries, stopping powers, coincidence summing corrections, and so on. The reader may find the discussions in Refs. \cite{Ili08,Row02} illuminating.

For many resonances we use the evaluated energies of Endt \cite{End90,End98}. However, for papers published after 1998 we are compelled to perform our own evaluation. Resonance energies can be measured directly using thick-target excitation functions or are derived by using measured excitation energies and the reaction Q-value ($E_r^{cm}=E_x-Q$; Sec. 5.1.1 in Paper I). We adopted in general the method that resulted in the smallest uncertainties. For reaction Q-values we use the evaluated results of Audi and collaborators \cite{Aud03}, unless more recently (after 2003) measured masses have been reported in the literature (see Tab. 1 of Paper II). 

Assignments of nuclear level spins and parities are especially precarious in the recent literature. As pointed out in the introduction to his 1990 evaluation \cite{End90}, Endt carefully distinguished between strong (that is, model-independent) and weak arguments when assigning $J^{\pi}$ values. Assignments based on weak arguments were placed in parenthesis by Refs. \cite{End90,End98}. In most papers published after 1998 this important distinction between strong and weak arguments is blurred and it requires now significant efforts by a reviewer to disentangle the arguments for $J^{\pi}$ assignments from different reaction and decay studies. We can hardly overstate to our colleagues the importance of strictly following the established rules for assigning spins and parities (see introduction of Ref. \cite{End90} and references therein). 

Regarding resonance strengths, we did not follow in general the procedure of the NACRE reaction rate evaluation \cite{Ang99}, where for a given resonance the strength is found from a weighted mean of values obtained in different measurements. Instead, whenever possible we normalize literature results to a set of carefully measured ``standard" resonance strengths (see Tab. 4.12 in Iliadis \cite{Ili07}). Note that no standard resonance strengths exist at present, for example, for ($\alpha$,$\gamma$) reactions or for (p,$\gamma$) reactions on any of the Ne isotopes.

Partial widths can be derived from measured resonance strengths, mean lifetimes or spectroscopic factors. See Ref. \cite{Ili07} for details. We prefer, whenever feasible, to calculate reaction rates by numerical integration (see Eq. (1) of Paper I) using partial widths as input instead of computing them analytically using the resonance strength (see Eq. (10) of Paper I). The former procedure automatically accounts for the low- and high-energy tails of a resonance and makes any artificial corrections (for example, for ``nonresonant" S-factor tails) obsolete. In most cases involving short-lived targets, we compute the partial widths by using measured spectroscopic factors and mean lifetimes of the corresponding levels in the mirror nucleus. This method and its justification has been described in detail by Iliadis et al. \cite{Ili99}. 

Uncertainties are treated in the following manner. For measured resonance energies the reported or derived mean value and the corresponding (``1$\sigma$") uncertainty is associated with the parameters $\mu$ and $\sigma$, respectively, of a Gaussian probability density function (Sec. 5.1.1 of Paper I). For measured resonance strengths or partial widths we associate the mean value and corresponding uncertainty with the expectation value and the square root of the variance, respectively, of a lognormal distribution (Sec. 5.1.2 of Paper I). The lognormal parameters $\mu$ and $\sigma$ are then computed from Eq. (27) of Paper I. If uncertainties are not available from the literature, we use certain global values to the best of our judgement: (i) the direct capture reaction rate is usually calculated using the method described in Secs. 2.1 and 5.1.3 of Paper I. It is sometimes assumed that this procedure represents a purely theoretical approach. However, this assumption is incorrect since the absolute magnitude of the direct capture S-factor is determined by using {\it experimental} spectroscopic factors in Eq. (35) of Paper I. We adopt in most of these cases an uncertainty (square root of the variance) of 40\% for the direct capture S-factor (see Eq. (5) of Paper I). This value is based on a systematic comparison of experimental spectroscopic factors from direct capture and from transfer reaction studies \cite{Ili04}; (ii) when particle and $\gamma$-ray partial widths are calculated from measured spectroscopic factors and $\gamma$-ray transition strengths, we assume uncertainties of 40\% and 50\%, respectively. The choice of these values is supported by a systematic comparison of partial widths \cite{Ili99} and by an uncertainty analysis of measured spectroscopic factors \cite{Tho99}; (iii) in exceptional cases we adopt spectroscopic factors and $\gamma$-ray transition strengths from the nuclear shell model. For the sake of consistency, we assume values of 40\% and 50\% for the uncertainties of shell model based particle and $\gamma$-ray partial widths, respectively.

Upper limits of particle partial widths are sampled according to the procedure outlined in Sec. 5.2.1 of Paper I. Specifically, the Porter-Thomas distribution of dimensionless reduced widths is obtained with mean values of $\langle \theta^2_p \rangle=0.0045$ for protons (or neutrons) and $\langle \theta^2_{\alpha} \rangle=0.010$ for $\alpha$-particles, and the distribution is truncated at the experimental upper limit of the dimensionless reduced width (see Eq. (38) of Paper I). If the spin and parity of a particular nuclear level is unknown, we assume formation of the expected (but yet undetected) resonance via s-waves ($\ell=0$).

Table 1 of Paper II contains a list of references for each reaction evaluated in the present work. The list is not exhaustive by any means, but provides the reader with an impression on the most recent or relevant work considered here. We hope that making our recommended input data available to the community represents a step forward in terms of the reproducibility and transparency of evaluated reaction rates. 
 
\section{Words of caution}

The input data presented in App. \ref{sec:nuclinput} are based on the most reliable information that we are able to extract from the published literature. They reflect our best current knowledge of these parameters. By no means can we exclude the possibility that, for example, a reported resonance strength was derived using the wrong stoichiometry, or that an incorrect $J^{\pi}$ value has been reported for a particular nuclear level. This issue should be kept in mind when drawing conclusions from our Monte Carlo reaction rate uncertainties.

Many test runs have been performed to ensure proper functioning of our new Monte Carlo code \texttt{RatesMC}. It is gratifying to see that the present reaction rates agree with those calculated using the previous code \texttt{RateErrors} \cite{Tho99} in those simple and restricted circumstances where the latter code provides accurate results (see discussion in Sec. 3.3 of Paper I). However, a number of issues are disregarded by \texttt{RatesMC} although their implementation is straightforward from a computational point of view: (i)  when integrating the rate contribution of a resonance numerically, and the resonance can be formed via two orbital angular momenta, $\ell$ and $\ell+2$, we only take the dominant contribution into account when scaling the particle partial width with energy (see Eq. (16) of Paper I); (ii) when integrating the rate contribution of a capture resonance numerically, and the $\gamma$-ray decay strength is fragmented, we only take the strongest primary transition (and the associated final state excitation energy) into account when scaling the $\gamma$-ray partial width with energy (see Eq. (17) of Paper I); (iii) when upper limits of partial widths are involved, the code samples over a Porter-Thomas distribution of dimensionless reduced widths, with a mean value of $\langle \theta^2_i \rangle$. The mean value is found from a least-squares fit, as discussed in Sec. 5.2.1 of Paper I, but the {\it uncertainty of the mean value} is not taken into account in the present version of the code; (iv) in some cases (i.e., for proton-induced reactions on $^{18}$F, $^{19}$Ne, $^{23}$Mg, $^{24}$Al, $^{29}$P, $^{32}$Cl and $^{39}$Ca), where information had to be extracted from the mirror nucleus, the analog assignments are not unambiguous. In principle, one could sample over a discrete distribution representing the different choices of analog assignments, but the present version of the code disregards this option; (v) for pairs of competing reactions, such as (p,$\gamma$) and (p,$\alpha$), or ($\alpha$,$\gamma$) and ($\alpha$,n), the partial widths entering in the rate calculations are correlated. Therefore, if our reaction rates are used in Monte Carlo nucleosynthesis studies, or are employed to derive the branching ratio (for example, 
$N_A\left<\sigma v\right>_{p\alpha}/N_A\left<\sigma v\right>_{p\gamma}$), then the resulting uncertainties on element abundances or branching ratios are likely overestimated. We did not implement the above options in the code since it is our intention to keep the formalism as simple and transparent as possible. Although we doubt that these issues will have major consequences, we may consider them in a future version of \texttt{RatesMC}.

The present Monte Carlo reaction rates also refine the motivation for future measurements in nuclear astrophysics. For example, if an experiment reduces the upper limit of a spectroscopic factor by an order of magnitude, then the ``upper limit" of the partial {\it classical} reaction rate is also reduced by that same factor. However, the Monte Carlo reaction rates, which are more reliable, behave in an entirely different manner, as explained at length in Sec. 4.4 of Paper II. Furthermore, there will be now a new emphasis on precise measurements of resonance energies since the associated uncertainties often give rise to long tails in the reaction rate probability density function (Secs. 4.2 and 4.3 of Paper II). 

Finally we would like to comment on ``direct" versus ``indirect" measurements. The expression ``direct" refers to the measurement of a reaction cross section of astrophysical interest. An ``indirect" measurement, on the other hand, refers to a study of some nuclear quantity (by using a reaction that is not necessarily the same as the one that occurs in the stellar plasma) from which the cross section of astrophysical interest may be partially inferred. For example, measurements of the reactions $^{21}$Na(p,p)$^{21}$Na, $^{24}$Mg(p,t)$^{22}$Mg or $^{24}$Mg($^{3}$He,n$\gamma$)$^{22}$Mg in order to study astrophysically important $^{22}$Mg states represent indirect studies of the $^{21}$Na(p,$\gamma$)$^{22}$Mg reaction, whereas an experiment using a radioactive $^{21}$Na beam on a hydrogen target is called a direct measurement of the $^{21}$Na(p,$\gamma$)$^{22}$Mg reaction\footnote{The reader should not confuse the expression ``direct measurement" with the unrelated term ``direct reaction" that denotes a single-step nuclear reaction.}. The calculation of reaction rates using results from indirect measurements necessarily involves the application of some nuclear model. Unfortunately, the systematic uncertainties introduced by a model are frequently difficult to quantify. For this reason it should be obvious that {\it a direct measurement is generally preferred over an indirect one, even if the estimated reaction rate uncertainties calculated using results from indirect measurements are relatively small.}

\section{Summary}
In the present paper of the series (Paper III) we publish the nuclear physics input data used to compute the Monte Carlo reaction rates presented in  Paper II. The reaction rates are calculated using the new method discussed in Paper I. Our input data, listed in the appendix, are based on the most reliable information that we are able to extract from the published literature. By making our recommended input data available to the community we intend to improve the reproducibility and transparency of the evaluated reaction rates. The reaction rate uncertainties given in Paper II are statistical in nature and do not account for unknown systematic errors in the nuclear input data listed here. The survey of literature for this review was concluded in November 2009. 

\section{Acknowledgement}
We would like to thank Ryan Fitzgerald for helpful comments. This work was supported in part by the U.S. Department of Energy under Contract No. DE-FG02-97ER41041. 

\clearpage
\appendix

\section{Explanation of tables} \label{sec:inptfiles}
As an example, we provide below a table with the nuclear data input used to calculate the reaction rates for a hypothetical reaction X(p,$\alpha$)Y with the code \texttt{RatesMC}. None of the entries in this table represent physical values, but are listed here for illustrative purposes only. All kinematic quantities are given in the center of mass reference frame. A row is disregarded as input if it begins with the symbol `!'. The meaning of each input row will be briefly explained. \\
\\
Sample nuclear data input:\\
\footnotesize

\normalsize
\vspace{6mm}
Explanation of input:
\vspace{4mm}
\begin{tabbing}
AAAAAAAAA\= \kill
Row 01:				\> Reaction label. \\
Row 02: 				\> Separator. \\
Row 03:				\> Projectile charge. \\
Row 04:				\> Target charge. \\
Row 05:				\> Charge of {\it exit particle}; it refers to the channel other than the entrance \\
                                         \> and the $\gamma$-ray channel; =0 if only two channels are open and radiative \\
					\> capture is the only possible reaction. \\
Row 06:				\> Projectile mass in atomic mass units. \\
Row 07:				\> Target mass in atomic mass units. \\
Row 08:				\> Mass of exit particle in atomic mass units. \\
Row 09:				\> Projectile spin. \\
Row 10:                            \> Target spin. \\
Row 11:                            \> Spin of exit particle. \\			
Row 12:			         \> Separation energy of incident particle in keV. \\
Row 13:                            \> Separation energy of exit particle in keV. \\
Row 14:                            \> Radius parameter in fm, used for calculating penetration factor. \\
Row 15:                            \> Label of $\gamma$-ray channel; channel 1 refers to the incident particle, channel \\
        				         \> 2 to the emitted quantum, and channel 3 to the spectator quantum. \\
Row 16:                            \> Separator. \\				
Row 17:                            \> Minimum energy cutoff (in keV) for numerical integration of rates. \\
Row 18:                            \> Number of random samples. \\				
Row 19:                            \> Flag for temperature output; =0 outputs results at all temperatures. \\
Row 20:                             \> Separator. \\
Rows 21-22:                       \> Comments. \\				
Row 23:                            \> Input for nonresonant contribution; \texttt{S,S',S''} are the parameters $S(0)$,\\
                                         \>  $S^{\prime}(0)$, $S^{\prime\prime}(0)$ of the astrophysical S-factor (Eq. 5 of Paper I) in units  \\
                                         \> of keVb, b, b/keV, respectively; \texttt{fracErr} is the fractional uncertainty, \\
                                         \> $\sqrt{V[x]}/E[x]$, of the effective S-factor (Eq. 8 of Paper I); \texttt{Cutoff Energy}  \\
                                         \> labels the energy $E_{\mathrm{cutoff}}$ (in keV) at which the S-factor is cut off at   \\
                                         \> higher energies; it is related to the cutoff temperature (see Eq. 9 of  \\
                                         \> Paper I) by the expression \\
\end{tabbing}
\vspace{3mm}
\begin{equation}
T_{9,\mathrm{cutoff}} = 19.92~E_{\mathrm{cutoff}}^{3/2}/\sqrt{Z_0^2Z_1^2\frac{M_0M_1}{M_0+M_1}}\notag
\end{equation}
\vspace{3mm}
\begin{tabbing}
AAAAAAAAA\= \kill                                      
                                         \> where $E_{\mathrm{cutoff}}$ is in MeV and all other quantities have the same meaning \\
                                         \>  as in Sec. 2 of Paper I. \\
Row 24:                            \> Input for a second nonresonant contribution, if needed. \\
Row 25:                            \> Separator. \\
Rows 26-29:                       \> Comments. \\
Row 30:                            \> Input for resonance contribution, one input row for each resonance; \\
                                         \> \texttt{Ecm,DEcm}: resonance energy and $1\sigma$ uncertainty; \texttt{wg,Dwg}: resonance \\
                                         \>  strength, $\omega\gamma$, and associated uncertainty; \texttt{Jr}: resonance spin; \\
                                         \> \texttt{G1,DG1,L1}: incident particle partial width, uncertainty, orbital angular  \\
                                         \> momentum quantum number; for a subthreshold resonance ($E_r<0$)  \\
                                         \> the dimensionless reduced width (Eq. 14 of Paper I) is listed instead \\
                                         \> of the entrance channel partial width; \texttt{G2,DG2,L2}: partial width,  \\
                                         \> uncertainty, angular momentum quantum number (or multipolarity for \\
                                         \> $\gamma$-rays) of emitted quantum; \texttt{G3,DG3,L3}: partial width, uncertainty, \\
                                         \> angular momentum quantum number (or multipolarity for $\gamma$-rays) of \\
                                         \> spectator quantum; \texttt{Exf}: excitation energy of level in residual nucleus \\
                                         \>  that is populated in primary transition; \texttt{Int}: =0 for analytical rate  \\
                                         \> calculation; =1 for numerical rate calculation (see Eq. 1 of Paper I); for  \\
                                         \> $E_r<0$ the rate contribution is always computed numerically; when the \\
                                         \> resonance strength is entered the rate contribution is always computed \\
                                         \> analytically, regardless of the flag value; \texttt{Ecm,DEcm,Exf} are in units of \\
                                         \> keV, while resonance strengths and  partial widths are in eV. \\
Row 31:                            \> Input for second resonance; in this example, the rate contribution is \\
                                         \> calculated from partial widths and the rate is chosen to be integrated \\
                                         \> numerically. \\
Row 32:                            \> Input for third resonance; in this example, the rate contribution is \\
                                         \> calculated from the resonance strength and the rate is necessarily \\
                                         \> computed analytically. \\
Row 33:                            \> Separator. \\
Rows 34-37:                      \> Comments. \\                                      
Row 38:                            \> Input for resonance contribution when only a partial (or reduced) width \\
                                         \> upper limit is available for at least one reaction channel; the number of \\
                                         \> upper limit channels must be less than the number of open channels; \\
                                         \>  the meaning of the input quantities is the same as for row 30, except \\
                                         \> that (i) resonance strengths are not allowed as input, and (ii) the mean \\
                                         \> value, \texttt{PT}, for the Porter-Thomas distribution of dimensionless reduced \\
                                         \> widths is entered for each upper limit channel (Sec. 5.2.1 of Paper I); \\
                                         \> upper limit channels are identified by a non-zero value for the partial \\
                                         \> (or reduced) width and a zero value for the corresponding uncertainty; \\
                                         \> in this example, an upper limit for the dimensionless reduced proton \\
                                         \> width is entered (since $E_r<0$). \\
Row 39:                            \> Input for second resonance; in this example, an upper limit for the \\
                                         \> dimensionless reduced $\alpha$-particle width is entered. \\
Row 40:                            \> Separator. \\
Rows 41-43:                      \> Comments. \\                                      
Row 44:                            \> Flag for interference between resonances of same spin and parity (Sec.  \\
                                         \> 2.4 of Paper I); \texttt{+}: positive interference; \texttt{-}: negative interference; \texttt{+-}:  \\
                                         \> unknown interference sign (a binary probability density function is then\\
                                         \> used for the random sampling; Sec. 4.4 of Paper I). \\ 
Rows 45-46:                       \> Input for two interfering resonances; the meaning of the input quantities \\
                                         \> is the same as for row 38, except that the rate contribution is always \\
                                         \> computed numerically. \\
Row 47:                            \> Separator. \\
Rows 48-49:                       \> Comments; references given in this section are summarized again after \\
                                         \> each table. The symbols \texttt{Gp}, \texttt{Ga}, \texttt{Gn}, \texttt{Gg} and \texttt{G} refer to the proton,  \\
                                         \> $\alpha$-particle, neutron, $\gamma$-ray partial width and total width, respectively; \\
                                         \> \texttt{Er} denotes the resonance energy in the center of mass system. \\
\end{tabbing}
%

%
%
\clearpage
\section{Nuclear physics input} \label{sec:nuclinput}
\footnotesize

\normalsize
\vspace{5mm}
References: Ajzenberg-Selove \cite{Ajz91}; Audi et al. \cite{Aud03}; Bartholomew et al. \cite{BA55}; Bommer et al. \cite{BO75}; Ferguson and Gove \cite{FE59}; French et al. \cite{FR61}; G\"orres et al. \cite{GO90}; Kuan et al. \cite{KU71,KU76}; Niecke et al. \cite{NI77}; Ramirez et al. \cite{RA72}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Audi et al. \cite{Aud03}; Cunsolo et al. \cite{CU81}; Gai et al. \cite{GA87}; G\"orres et al. \cite{GO92}; Lugaro et al. \cite{LU04}; Tilley et al. \cite{Til95}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Becker et al. \cite{BE82}; G\"orres et al. \cite{Goe00}; Kieser et al. \cite{Kie79}; Rolfs, Berka and Azuma \cite{Rol73a}; Rolfs, Charlesworth and Azuma \cite{Rol73b}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Angulo et al. \cite{Ang99}; de Oliveira et al. \cite{OL96}; Wilmes et al. \cite{WI02}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Davids et al. \cite{Dav03b}; Dufour and Descouvemont \cite{Duf00}; Goriely, Hilaire and Koning \cite{Gor08}; Kanungo et al. \cite{Kan06}; Langanke et al. \cite{Lan86}; Mao, Fortune and Lacaze \cite{Mao95}; Mythili et al. \cite{Myt08}; de Oliveira et al. \cite{OL96}; Tan et al. \cite{Tan05}; Tilley \cite{Til95}.   
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Ajzenberg-Selove \cite{Ajz72}; Angulo et al. \cite{Ang99}; MacArthur et al. \cite{Mac80}; Mao, Fortune and Lacaze \cite{Mao96}; Mayer \cite{May01}; Mohr \cite{Moh05}; Tilley et al. \cite{Til98}.   
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Angulo et al. \cite{Ang99}; Champagne and Pitt \cite{CH86}; La Cognata et al. \cite{La08}; Lorentz-Wirzba et al. \cite{LO79}; Tilley et al. \cite{Til95}; Wiescher et al. \cite{WI80}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Almanza et al. \cite{AL75}; Lorentz-Wirzba et al. \cite{LO79}; Mak et al. \cite{MA78}; Murillo et al. \cite{MU79}; Orihara, Rudolf and Gorodetzky \cite{OR73}; Sellin, Newson and Bilpuch \cite{SE69}; Yagi \cite{YA62}. 
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Berg and Wiehard \cite{Ber79}; Buchmann, D'Auria and McCorquodale \cite{Buc88}; Chouraqui et al. \cite{Cho70}; Dababneh et al. \cite{Dab03}; Descouvemont \cite{Des88}; Endt \cite{End98}; Giesen et al. \cite{Gie94}; Goldberg et al. \cite{Gol04}; Graff et al. \cite{Gra68}; Trautvetter et al. \cite{Tra78}; Vogelaar et al. \cite{VO90}. 
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Nesaraja et al. \cite{NE07}; Utku et al. \cite{UT98}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Adekola \cite{AD09}; Bardayan et al. \cite{BA01}; Bardayan et al. \cite{BA02}; Bardayan et al. \cite{BA04}; Bardayan et al. \cite{BA05}; Chae et al. \cite{CH06}; Nesaraja et al. \cite{NE07}; Utku et al. \cite{UT98}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Audi et al. \cite{Aud03}; Coszach et al. \cite{CO94}; Tilley et al. \cite{Til98}; Vancraeynest et al. \cite{VA98}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Anttila et al. \cite{antilla}; Endt and van der Leun \cite{End78}; Firestone \cite{endsf}; Iliadis \cite{ili2}; Iliadis et al. \cite{Ili08}; Mukherjee et al. \cite{muk2}; Rolfs et al. \cite{Rol75}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Anantaraman et al. \cite{raman}; Endt and van der Leun \cite{End78}; Fifield et al. \cite{fif}; Firestone \cite{endsf2}; Highland and Thwaites \cite{high}; Schmalbrock et al. \cite{sch}; Smulders \cite{smu}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Endt \cite{End90}; G\"orres et al. \cite{Gor83}; Hale et al. \cite{Hal01}; Powers et al. \cite{Pow71}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Longland et al. \cite{Lon09}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Longland et al. \cite{Lon09}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: He et al. \cite{He07}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Endt \cite{End90}; Lotay et al. \cite{Lot08}; Tomandl et al. \cite{Tom04}; Visser et al. \cite{Vis07,Vis08}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Cseh et al. \cite{cseh}; Draayer et al. \cite{dr74}; Endt \cite{End90}; Firestone \cite{endsf2}; Lyons et al. \cite{ly69}; Maas et al. \cite{maas}; Smulders and Endt \cite{smE}; Strandberg et al. \cite{str08}; Tanabe et al. \cite{ta83}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Endt \cite{End90}; Endt and Rolfs \cite{EnR87}; Iliadis \cite{Ili89}; Iliadis et al. \cite{Ili96}; Iliadis et al. \cite{Ili01}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Endt and Rolfs \cite{EnR87}; Iliadis \cite{Ili89}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Endt and Rolfs \cite{EnR87}; Iliadis \cite{Ili89}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Champagne et al. \cite{Cha90}; Iliadis et al. \cite{Ili90}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Herndl et al. \cite{Her95}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Endt \cite{End90}; Iliadis et al. \cite{Ili96}; Parpottas et al. \cite{Par04}; Wrede \cite{Wre09b}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Lotay et al. \cite{Lot09}; Vogelaar \cite{Vog89}; Vogelaar et al. \cite{Vog96}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Chronidou et al. \cite{Chr99}; Endt \cite{End98}; Harissopulos et al. \cite{Har00}; Iliadis et al. \cite{Ili01}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Champagne et al. \cite{Cha88}; Endt \cite{End98}; Endt and Booten \cite{EnB93}; Iliadis et al. \cite{Ili01}; Timmermann et al. \cite{Tim88}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Moon et al. \cite{Moo05}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Endt \cite{End98}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Endt \cite{End90}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Bardayan et al. \cite{Bar07}; Endt \cite{End90}; Iliadis et al. \cite{Ili01}; Makh et al \cite{Mac73}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Endt \cite{End98}; Iliadis et al. \cite{Ili93}; Kalifa et al. \cite{Kal78}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Endt \cite{End98}; Kalifa et al. \cite{Kal78}; MacArthur et al. \cite{Mac85}; Ross et al. \cite{Ros95}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Axelsson et al. \cite{Axe98}; Fynbo et al. \cite{Fyn00}; Wrede et al. \cite{Wre09}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Endt and van der Leun \cite{End78}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Herndl et al. \cite{Her95}; Schatz et al. \cite{Sch05}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Johnson et al. \cite{Joh74}; Endt and van der Leun \cite{End78}; Iliadis et al. \cite{Ili94}; Iliadis et al. \cite{Ili01}; R\"opke et al. \cite{Roe02}; Ross et al \cite{Ros95}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Bosnjakovic et al. \cite{Bos68}; Endt and van der Leun \cite{End78}; Endt \cite{End98}; Iliadis et al. \cite{Ili94}; R\"opke et al. \cite{Roe02}; Ross et al \cite{Ros95}.
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
References: Trinder et al. \cite{Tri99}; Yazidjian et al. \cite{Yaz07}. 
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
\clearpage
\footnotesize

\normalsize
\vspace{5mm} 
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
\clearpage
\footnotesize

\normalsize
\vspace{5mm}
\clearpage
\footnotesize

\normalsize
\vspace{5mm}

\end{document}